Q.Two pistons of hydraulic press have diameter of 30cm and 2.5cm.


Find the force exerted by the longer piston when 50kg wt is placed on smaller piston.If the stroke of the smaller piston is 4cm, find the distance through which the longer piston would move after 10 strokes

Grade:12th pass

2 Answers

Vijay Luxmi Askiitiansexpert
13 years ago

Hi,

 

Please don't post the same question twice..

Chetan Mandayam Nayakar
13 years ago

ratio of forces =(30/2.5)2=144, force exerted by larger piston=50*144=7200kg wt

larger piston will move a distance= 4*10/144=5/18=0.27cm